OpenStudy (anonymous):

what is 37.5% of 16 what is 0.25% of 900?

OpenStudy (anonymous):

Need answer ASAP due in an hour

zepdrix (zepdrix):

25% can be written as \(\dfrac{25}{100}\quad=\quad \dfrac{1}{4}\) But we don't want 25%, we want .25% or one-hundredth of 25%, \(\dfrac{1}{4}\cdot\dfrac{1}{100}\quad=\quad\dfrac{1}{400}\) To find 0.25% of 900 we will multiply this percent (now in fraction form) by 900. \[\Large \frac{1}{400}\cdot900\quad=\quad?\]

OpenStudy (anonymous):

I dont need HOW to solve it I need the direct answer

zepdrix (zepdrix):

That's not what this website is for, go cheat somewhere else.